Journalist Report – January 10th

Journalist Report

Nicole Chan, Crew 290 Journalist


Sol 3

At the end of Sol 2, we successfully carried out our callsign ceremony, officially naming Melon (Madelyn), Chopper (Rebecca), Freebee (Anja), Funk (Ben), Roots (Anna), and PODO (Nicole). After two busy Sols of onboarding and settling into our rotation, the crew was granted a later start in their work day. We all awoke from a great night of sleep to the smell of coffee and bread, courtesy of Anja (Freebee) testing out the bread machine before bed. The sun was already out and shining through our Hab windows and made for a pleasant start to Sol 3. The crew seized the day and went straight to work after a breakfast of pancakes from Freebee (stepping up her cooking for sure). With chocolate chips and plenty of condiments (peanut butter and Nutella), it was a unanimously amazing meal and brings back great memories of having similar pancakes on Earth.

In preparation for EVA 4, Freebee worked on troubleshooting any issues with REMI and was successful in the setup. Ben (Funk) and Anna (Roots) were performing soil salinity tests in the GreenHab while Rebecca (Chopper) continued setting up the ScienceDome for the sample collection EVA and Madelyn (Melon) evaluated the area’s geology for future EVA sites. Our lunch was a nostalgic PB&J sandwich with Freebee’s homemade bread. Though dense, the bread had a mild sweetness to it and gave us the energy to continue with our Sol 3 objectives. Shortly after, EVA 4 crew, which consisted of Freebee, Funk, and myself (Podo), began suiting up and headed out to the Hab’s surrounding area with REMI. It was a very productive and successful EVA! REMI roamed around several geological points of interest and collected radar data while Funk and Podo identified viable sample collection sites and proceeded with temperature and salinity tests of the soil alongside sample collection. Funk settled into a routine where he was responsible for digging at the sites while I took pictures of the site as well as the measurements of the soil sample depth. The shovel we had on hand took a toll as Funk was determined to collect samples on the surface in smectite, at 3 inches deep in clay, and at 6 inches deep into shale at the site. The physical demand of the task was well worth the scientific output of the samples. We are all excited to see what we will find through the soil tests. With Melon on CAPCOM, we made plenty of field notes that will aid in our analysis.

After returning from the EVA, Funk headed over to the GreenHab to mist the precious plant babies and performed valuable knowledge transfer with Crew 289’s GreenHab officer Riya to understand how to better support the plant growth. In addition, the plants had a more diverse music therapy today so hopefully they feel loved and cared for by our crew. Chopper was hard at work in the ScienceDome at the laminar fume hood to set up microfluidic device tests to process the soil samples from EVA 4. Dinner is served by Funk tonight and smells amazing as I am writing this report. It will be Hearty Chicken and Potatoes with a side of mixed vegetables and a sour cream dill sauce (from our dill harvest yester-sol). We cannot wait to dig in and look ahead at the rest of this week’s EVAs and review science objectives together. The science experiments are all set up and as Chopper introduces the crew to the testing protocol, we look forward to making awesome findings in the coming days.

Photo of the Day: 290-01_10_2024 eva 4 begins

Operations Report – January 10th

Crew 290 Operations Report 10-01-2023

SOL: 3

Name of person filing report: Anna Tretiakova

Non-nominal systems: Toilet Tank Sensor

Notes on non-nominal systems: The toilet tank was at full on the sensor and was emptied. After hearing that everything went through the pipe, the sensor still showed that the tank is ⅔ full. Mission support was contacted and they said that is an issue with some gunk stuck on the sensor. In addition to the sensor problems, the whole room stinks. Moving forward, is it possible to get a febreeze? Hopefully the emptying of the septic tank solves this issue.


Spirit rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Opportunity rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Curiosity rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

Perseverance rover used: No

Hours: (Before EVA)

Beginning Charge: (Before EVA)

Ending Charge: (After EVA and before recharging)

Currently Charging:

General Notes on Rovers: Rear view mirrors will be checked and tightened during EVA 5

Summary of Hab Operations:

WATER USE: 7.1 gallons

Water (static tank): 412.4 gallons

Static tank pipe heater (on or off): on

Static tank heater (on or off): on

Toilet tank emptied: yes

Summary of Internet: internet is nominal during times 18:00 to 22:00

Summary of Suits and Radios: Radios nominal. Suits nominal. Suit 8 had a blown fuse, it was replaced and now suit 8 is nominal.

Summary of Greenhab Operations:

Plants were misted 3 times and the salinity of a few plants was tested. More details in the Greenhab report.

WATER USE: 5.5 gallons

Heater: on

Supplemental Light: off

Harvest: Not yet.

Summary of Science Dome Operations:

Set up the Science dome for MADMEN experiments. Cleaned and sterilized the hood. Started the control test: so far everything is going as planned. The incubator is operational. Optimized the microscope to be able to see the surface of the microtubule devices using two optical lenses (4x and 10x). The day was finished with set up and preliminary tests of the ATP sensor.

Dual Split Heater or AC: On

Summary of RAM Operations: (Tools used, work done):

Crew scientists continued working on their robot in the RAM. The RAM was also used as the airlock for EVA 4. Measuring tape was used to measure the water level.

Summary of any observatory issues: Nominal

Summary of health and safety issues: All crew in good health and happy spirits. No safety issues.

Questions, concerns and requests to Mission Support:

None at the moment.

GreenHab Report – January 10th

Crew 290 GreenHab Report 10-01-2024 (Sol 3)

Greenhab Officer: Ben Kazimer, Callsign “Funk”

Heater on, fan off, door closed

Average temperatures: 0955: 61.2⁰F & 22%, 1222: 77.2⁰F & 22%, 1721: 69.6⁰F & 51%

Hours of supplemental light: Nominal 4 hours overnight

Daily water usage for crops: 5 gallons

Daily water usage for research and/or other purposes: 0.5 gallons

Water in Blue Tank 176.3 gallons

Time(s) of watering for crops: 1133: 3 gallon watering and 32 oz misting, 1705: 1 gallon watering and 32 oz misting

Changes to crops: pruned some dead flowers off the cucumbers in the large blue planter. Added two Earth-based potato chunks on toothpicks to beakers with water to promote sprouting.


I was able to spend a good amount of time in the greenhab today, it was fantastic! I checked in at 0955 to see a relatively cold environment of 61.2⁰F. I had a chat with Mission Support on nominal greenhab status, to which he sent me some words of wisdom from Crew 289’s GHO (Thanks Riya 😊). Going forward, I’m going to mist the green guys a few times while maintaining 3-5 gallons of direct watering. For most of the morning, Engineer Anna and I measured soil salinity levels of many plants to have as reference for any future changes (fertilizer, repotting, etc.). We were able to setup a makeshift laboratory in the greenhab with my plant-viewing benches and it worked quite well! I even found a tiny pot full of unused potting soil that I plan to use for Earth-grown potato sproutings. I suspended two potato chunks for sprouting over separate beakers full of water, fingers crossed we see some roots! Last note is the now regular music rundown of the day: I downloaded the highlight release of Live in Europe ’72 from the GD and I certainly enjoyed it so hopefully the plants did too. I was apprehensive to expose our crops to rap music, but I felt like a little switch up during the pre-dinner checkup wouldn’t hurt… Sol 4 will tell us all if the greenhouse inhabitants are welcoming of OVO music.

Harvest: N/A

Support/supplies needed: None!

EVA Report – January 10th

Crew 290 EVA Report 10-01-2024

EVA # 04

Author: Madelyn Hoying

Purpose of EVA: Confirm all systems on REMI robot are nominal after shipping and reconstruction, and to begin collecting initial data from at least 2 different terrain types.

Start time: 1400

End time: 1600

Narrative: This was a repeat of the original plan for EVA 03 to confirm all systems on REMI robot are nominal and begin sampling for MADMEN. REMI sampled six different sites within 180 meters of the Hab. The focus of the sites was mostly on sandy and gravelly terrain near small streamways. All systems were nominal during the EVA, and the crew had little to no difficulty running commands and controlling the robot with suits on. A new setup procedure in the RAM airlock also helped alleviate some struggles from EVA 03. For MADMEN, the use of the salinity meter in the field was confirmed without obstruction by the suits or gloves. Regolith and rock samples were collected and easily identified as samples of interest by the field scientists and CAPCOM, with temperature and additional data collected at each sample depth. Overall, this was a resounding success!

Destination: within 200m of MDRS

Coordinates (use UTM WGS 84): N518169, E4250939

Participants: Anja Sheppard (Crew Scientist), Ben Kazimer (GreenHab), and Nicole Chan (Crew Journalist)

Road(s) and routes per MDRS Map: Exit the RAM airlock with REMI and walk to sites of interest within a 200m radius

Mode of travel: Walking

Vehicles to be used (If applicable): None

Sol Summary – January 10th

Crew 290 Sol Summary Report 10-01-2024

Sol: 3

Summary Title: Hitting rock bottom…in a good way

Author’s name: Madelyn Hoying

Mission Status: Nominal

Sol Activity Summary:

Things are starting to come together! After a breakfast of chocolate chip pancakes and homemade bread courtesy of Anja (callsign: Freebee), the crew kicked off the day with science prep, GreenHab upkeep, and habitat housekeeping. After a light lunch, EVA 04 prep commenced for Freebee, Ben (Funk), and Nicole (PODO), with Madelyn (Melon) as CAPCOM. On EVA 04, REMI and Freebee sampled six different sites within 180 meters of the Hab. The focus of the sites was mostly on sandy and gravelly terrain near small streamways. All systems were nominal during the EVA, and the crew had little to no difficulty running commands and controlling the robot with suits on. A new setup procedure in the RAM airlock also helped alleviate some struggles from EVA 03. EVA 04 was also the first sample collection for MADMEN. Funk performed human factors assessments of salinity field tests and collected the necessary samples from shale in the Curtis foundation near the Hab. EVA 04 was a success and provided Rebecca (Chopper) with samples to run to begin the microbiological analysis goals of Project MADMEN. As always, Anna (Roots) proved to be an excellent Crew Engineer and managed to troubleshoot multiple things as they came up, preventing progression towards serious issues. The crew is settling in for a dinner made by Funk and looking forward to some more crew bonding activities tonight!

Look Ahead Plan: Engineering EVA, science planning

Anomalies in work: pH meter still not functional

Weather: Warm on EVA!

Crew Physical Status: nominal

EVA: 04 Science tasks for REMI and sample collection, with Anja, Ben, and Nicole

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA Report, EVA Request

Support Requested:

Septic tank and toilet solution (see ops report)
For engineering EVA to fix the RAM ramp: any updates on use of a shovel and where to take regolith from?
Can we use GreenHab shovels to assist in sample collection on EVA?

GreenHab Report – January 13th

Crew 290 GreenHab Report 13-01-2024 (Sol 6)

Greenhab Officer: Ben Kazimer

Heater on, fan off, door closed

Average temperatures: 0925: 71.2⁰F, 17%, 1729: 75.4⁰F, 20%

Hours of supplemental light: 4.5 (4 overnight and ½ from misting after sunset)

Daily water usage for crops: 5.25 gallons

Daily water usage for research and/or other purposes: None!

Water in Blue Tank 158.05 gallons

Time(s) of watering for crops:

0925: 3 gallons watering, ½ gallon misting, 1729: 1.5 gallon watering, ¼ gal misting

Changes to crops: Harvested a fair bit a greens for salad tonight, paid extra attention to moisture levels in tomato plant soil


Today was a bit lighter on Greenhab operations due to preparing, executing, and unloading from a 4-hour EVA to Candor Chasma (see other crew member reports for details!) I was still able to get all the watering and misting in, and I even swapped the environmental sensors between the Science Dome and Greenhab. Chopper had been harboring a more capable sensor that displays the current status, as well as the 24-hour max and min reading, so I took that and gave her my plain sensor. I cleaned up some dead pieces of plant throughout the growing areas and made sure that all sprouting plants were not choking out their neighbors (the pea shoots put up a good fight against the cilantro). In Tomatoes news, Tomato number 4 now has 2 yellowing fruit! I’m hopeful Crew 290 will experience some ripe tomatoes by the end of our mission (291 is going to have enough to make a whole pot of Sunday sauce!). Tomorrow is our rest day so I’m looking forward to spending some time going through all the plants in detail to trim unhealthy branches and rearrange crops as needed. Lastly, if you have ever wondered how a plant’s growth rate would respond to about half an hour of Louis Prima, Dean Martin, and Frank Sinatra, then tune into my sol 7 report!

Harvest: 73 g Kale, 51 g Red Cabbage, 11 g Parsley

Support/supplies needed: Another idea for future greenhouse upgrades, a nice stereo system with surround sound would be cool. The plants seem to love the music as much as I do 😊

Journalist Report – January 12th

Journalist Report

Nicole Chan, Crew 290 Journalist


Sol 5

This morning, the crew made staggered entrances to the kitchen to enjoy some cookies and bread made by Anja (Freebee) last night. With today’s EVA 6 being our longest yet, our EVA crew, Madelyn (Melon), Anna (Roots), and Freebee began preparing for their adventure to Vallaris Marineris. This region has promising geological features that Melon hoped to target, from evidence of river flow in the area. The rest of the crew did our best to help feed them (with Freebee’s tasty risotto from last night) and suit up before departing from the Hab at 1255. The goal of this EVA is to collect samples from two sites of interest for MADMEN science objectives while Freebee is to drive REMI around those sites to collect more radar data.

At the base, Ben (Funk) was responsible for CAPCOM while Rebecca provided scientific field test support for the EVA crew. I (PODO) experimented with the pH probe to determine the electrical charge buildup that is necessary for accurate pH measurements on field tests. Currently, its metal panels require 5 minutes of charge buildup to get an accurate reading of the buffers we have on hand. Although this was not the expected measurement protocol, it has been a great workaround for the pH meter issues we had during the initial setup period.

The EVA crew ended up collecting samples at the mouth of the Tharsis Montes riverbed. They were even able to find fossils! While they performed ATP field measurements and salinity tests of the soil, the crew also collected plenty of interesting rocks for analysis. The most worthy of note amongst those is the gypsum-resembling rock. It’ll be analyzed and properly categorized under the microscope in the coming days. Overall, it was a great success! We congratulated the EVA crew at their return and are super excited to “dig into” (ba-dum-tsh) the samples they collected. Today we also got the chance to image the microfluidic devices with the Sol 3 and found bacteria in them! Very exciting indeed. (To be fair, we’re all in a very excited mood right now particularly while we’re smelling the tasty Chana Aloo that Funk is cooking up.)

Looking ahead, we have another long EVA tomorrow that will target sites at Candor Chasma. With a rover ride included, we can’t wait to bring back some sick samples to the crew. We’ll keep you posted!

Photo of the Day: 290-01_12_2024 in da valley.JPG

Sol Summary – January 12th

Crew 290 Sol Summary Report 11-01-2024

Sol: 5

Summary Title: Signs of life

Author’s name: Madelyn Hoying

Mission Status: Nominal

Sol Activity Summary:

The day began with science, with Chopper (Rebecca), Roots (Anna), Freebee (Anja), and PODO (Nicole) working out the pH and microfluidics studies. The first signs of life were observed out of the microfluidic devices! Melon (Madelyn) laid out a checklist for sample site selection in the field to optimize EVA output. After a lunch of leftover risotto, the EVA crew of Melon, Freebee, and Roots, with Funk (Ben) on CAPCOM, ventured out for the furthest expedition yet with an ambitious goal to reach Lith Canyon. Unfortunately, the rover batteries reached their limit at about halfway, necessitating a turn-around point near Tharsis Montes. Two diverse sample sites were investigated, with one in a valley between two Morrison Formation hills and another in a dried riverbed heading east into the Tharsis Montes region. Field tests gave clear indications of microbial life at the riverbed site, so the crew is excited to begin lab testing and compare those results to the smectite regions! REMI had another solid day of data collection, managing to squeeze 10 sites into the EVA. Funk is cooking his own, specially-developed space food recipe tonight of chana aloo, and the crew is impatiently waiting for a taste test. With water usage still remarkably low post-shower, the crew is looking forward to showers after another long sample collection mission tomorrow.

Look Ahead Plan: Sample collection in the field and science!

Anomalies in work: none

Weather: Windy, but less so than Sol 4

Crew Physical Status: nominal

EVA: 05 engineering EVA, with Anja and Anna

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA Report, EVA Request

Support Requested:

Trash pickup tonight in the rear airlock
There is an old handheld GPS that was found on the shelf with the walkie talkies in the ground floor of the Hab. Is there a charger for this? It ran out of battery in the field today.

GreenHab Report – January 12th

Crew 290 GreenHab Report 12-01-2024 (Sol 5)

Greenhab Officer: Ben Kazimer

Heater on (Shifted to 65⁰F), fan off, door closed

Average temperatures: 0845: 70.7⁰F, 15%, 1158: 86.7⁰F, 10%, 1655: 68.7⁰F, 44%

Hours of supplemental light: Nominal 4 hours overnight

Daily water usage for crops: 5.25 gallons

Daily water usage for research and/or other purposes: None!

Water in Blue Tank 163.3 gallons

Time(s) of watering for crops:

0845: 3 gallons watering, ¾ gallon misting, 1158: ½ gallon misting, 1639: ½ gallon watering, ½ gal misting

Changes to crops: Rearranged positioning of tomatoes and moved them slightly away from the greenhab wall. Pruned all stressed/discoloring leaves on tomatoes and cucumbers


Today was a good sol in the world of the greenhouse. The plants got a healthy dose of water for breakfast and several mistings throughout the day. I noticed the tomatoes on the back of the main platform, closest to the outside wall, were slightly wilting and looking rather somber. I took matters into my own hands (even though I guess they’re the only hands available) and rearranged the positioning of all the tomato plants, to be further away from the wall and swapped the arrangement to alternate who’s getting the cold side of the bed (Tomato No. 9 stayed put mainly as I didn’t want to interrupt his flow and partially due to favorite child actions). If it’s not the cold, it could also be over exposure to the sun as they are closely in line with the sun’s trajectory. While the plants were doing the Martian Shuffle, I also trimmed off all the dead leaves. They received a good diversity of sonic vibrations today, the highlight being a double dose of Billy Strings! In preparation for the long-awaited Chana Aloo dinner, I was able to harvest some fresh cilantro for garnish. I clipped the herbs radially working inward as to prevent overlap with other herbs in the garden, the neighboring kale and dill appreciated this. I’m not sure if it’s because our Martian kitchen is devoid of fresh spices, but this cilantro tasted just as great as the arugula did yestersol. The rest of the crew shared this sentiment as the curry with garnish was a hit!

Harvest: 29 grams cilantro

Support/supplies needed: Just food for thought in terms of future greenhab upgrades, an automatic misting system or at least electrically driven handheld device would be great! It is tedious to deliver the misting (for managing humidity levels) with a spray bottle, but still manageable. Also, an upgrade to the environmental sensor would make a big difference, for more accurate data and possibly even a timelog. Knowing what’s going on overnight would be great!

Copyright © The Mars Society. All rights reserved. | Main Site